A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1870–1879

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

&c [233].

-and [40].

10th [352, 73, 48]. 14th [19]. 16th [214]. 17th [73, 160]. 1870 [19, 10]. 1871 [48, 36]. 1872 [63, 73]. 1874 [197, 163, 147]. 1875 [194, 197, 224, 214]. 1876 [276, 261, 262, 298]. 1877 [297]. 1879 [386, 385]. 18th [7, 385, 104, 386]. 19th [72, 342]. 1st [262].

20th [63, 352, 297]. 21st [19, 45, 261]. 22nd [36, 73]. 22nd/10th [73]. 23rd [10]. 25th [7, 330]. 26th [287]. 27th [251, 352]. 29th [73, 197]. 29th/17th [73].

30th [97, 142, 158, 143, 196, 212, 250, 259, 286, 295, 329, 375, 194]. 31st [10, 63, 97, 172, 143, 187, 228, 238, 270, 251, 281, 309, 287, 319, 330, 375, 36, 197].

5th [352].

7th [262].

9th [88].

Aberdeen [301]. Abroad [125]. Abstract [303]. Abstracts [152]. Accomodation [68]. According [205]. Account [375, 91]. Accounts [91]. Accumulations [311]. Accumulative [99]. Act [194, 58, 74]. Acts [26, 254]. Additions [109, 123, 142, 158, 172, 187, 196, 212, 228, 238, 250, 259, 270, 281, 286, 295, 309, 294, 319, 329, 339, 340, 350, 396, 383, 362, 374, 395]. Address [301, 385, 163, 263, 231, 45, 72, 105, 299, 104, 160, 201, 214, 261, 106, 344, 15, 264, 47, 48, 147, 346, 297, 342, 387]. Administration [132, 156]. Advancement [19, 7, 263, 105, 299, 201, 15, 47, 147]. Affected [174]. Affiliated [274]. Agricultural [83, 166, 134, 307, 347, 389]. Agriculture [233, 385, 25]. Algeria [255]. America [224]. American [25, 148]. Amherst [224]. among [145, 125]. Amount [80]. Analysis [378]. Anniversary [10, 63, 97, 144, 143, 198, 197, 252, 251, 288, 287, 331, 330, 376, 375, 36]. Annual [163, 231, 264, 391, 346, 254]. Answer [98]. any [313]. Aptitude [25]. Area [120, 205]. Arithmometer [343]. Armaments [133]. Arrangement [46]. Aspects [289]. Assets [91]. Association [19, 7, 163, 263, 231, 105, 299, 201, 15, 264, 47, 147, 346, 291, 300, 344, 48, 304, 387]. Attempt [98]. August [73, 147, 7, 88, 300, 299, 201, 344, 352, 47, 387]. Australasian [394, 116]. Australia [192, 226, 217]. Austria [312]. Average [380, 391].

Bank [194, 40, 312, 58, 391, 107, 74, 312]. Bank-Rate [40]. Banker [130, 245, 368]. Banking [316, 116, 230, 379, 80, 107, 80]. Banks [180, 392, 303, 391]. Baronetcies [248]. Basis [91]. be [191, 99, 91, 313]. Become [392]. been [313]. Belfast [147]. Belgium [204, 312]. Beneficial [175]. Bengal [120, 117, 54]. Between [360, 188, 272, 371]. Bibliography [327]. Bill [24, 140]. Bills [80]. Births [140, 302]. Boat [359]. Bonds [99]. Book [91, 246, 65, 203]. Books [308, 338, 318]. Booth [140]. both [80]. Bradford [105]. Branches [391]. Brigade [235, 277, 314, 357]. Brighton [231]. Bristol [201]. Britain [166, 16, 313, 35, 80]. British [300, 105, 299, 201, 344, 15, 47, 147, 304, 387, 19, 7, 247, 291, 263, 203, 125, 377]. Brussels [191]. Buda [276, 262, 298]. Buda-Pesth [276, 262, 298]. Bullion [230]. Bureau [203].

Caird [301]. Called [23]. Canada

[205, 345]. Canal [137, 65]. Canals [253]. Cap [88]. Capital [392, 91, 311]. Casualties [324, 369]. Causes [107]. Celtic [377]. Census [6, 205, 247, 370, 117]. Charged [107]. Charter [74]. Charts [276]. Cheltenham [346]. Children [265]. Chinese [317]. Chubb [89]. Church [4]. Circulation [230, 391, 80]. Claims [23]. Clauses [88]. Clearing [130, 131, 245, 368, 391]. Climate [233]. Co [136, 146]. Co-Operative [146, 136]. Coal [108]. Coinage [3]. Coinages [139]. Coincidence [360]. Collection [46]. College [197, 224, 251, 287, 330, 375, 160, 334]. Colonial [273]. Colonies [394, 55]. Colony [364, 217]. Commerce [223, 233, 272]. Commercial [100, 150, 127, 193, 244, 284, 323, 355, 367, 253, 313, 40]. Commissariat [348]. Commission [119, 335, 352]. Commodities [353]. Companies [253, 2]. Comparative [52, 302]. Compared [221, 177]. Compulsory [366]. Concerns [253]. Concluding [298]. Condensed [267]. Condition [167, 204, 161]. Conditions [313]. Conference [192]. Conferences [336]. Conferred [248]. Congress [352, 119, 191, 276, 335, 73, 163, 231, 106, 264, 262, 298, 352, 346, 215]. Connected [90]. Considered [289]. Consul [255]. Consul-General [255]. Consumption [222, 66, 174, 17]. Contagious [26, 254]. Continent [16, 98]. Convict [167]. Corn [380]. Cost [221, 283]. Cotton [17]. Council [10, 6, 63, 97, 143, 197, 251, 287, 330, 375, 36].

Counties [391]. Countries [322, 302, 321, 98, 371]. Country [79]. County [155]. Courts [115]. Cousins [188, 200]. Crises [40]. Crisis [58, 379, 11]. Critique [74]. Crop [210]. Crops [205, 390]. Cultivated [98]. Cultivation [17]. Currency [239, 31]. Current [91, 107]. Customs [151]. Cycles [359].

Dangers [93]. Day [94]. Death [149, 179, 360, 161, 199]. Death-Rate [179, 360]. Death-Rates [149, 161, 199]. Deaths [90, 140, 302, 125]. Debts [114, 322]. Decade [60]. December [197, 10, 63, 97, 172, 143, 228, 270, 251, 309, 287, 330, 375, 36]. Decreasing [253]. Deductions [391]. Deferred [253]. Delivered [385, 45, 72, 104, 160, 214, 261, 297, 342]. Demography [352]. Denmark [80]. Department [231, 48, 346, 163, 264]. Depend [313]. Depreciation [268]. Depression [332]. Deputation [140]. Derived [326]. Diagrams [276, 312]. Different [145, 272]. Discipline [54]. Discontinuance [304]. Discussion [81]. Diseases [26, 254]. Distribution [41]. Doctrines [300]. Domesday [246]. Doomsday [91]. Dr. [276]. Drains [11]. Dublin [344]. During [109, 123, 142, 158, 172, 187, 196, 212, 228, 182, 238, 250, 259, 270, 235, 281, 286, 295, 309, 277, 329, 314, 396, 190, 17, 218, 125, 40, 3, 319, 357, 391]. Duty [174]. Dwellings [175].

Each [205]. Early [225]. Economic [291, 263, 105, 299, 275, 201, 15, 47, 147, 304, 387]. Economical [349, 361, 373, 382]. Economy [162, 327, 231, 264, 48, 346, 163]. Edinburgh [47]. Education [366, 24]. Effect [175, 199]. Effective [154]. Effects [188, 379, 321, 40, 31]. Egypt [233]. Eighteen [203]. Eighteenth [163]. Eighth [63, 73, 344]. Elections [126]. Elementary [24]. Empire [67]. Endangered [313]. Ended [10, 63, 97, 142, 158, 172, 143, 187, 196, 212, 228, 197, 238, 250, 259, 270, 281, 286, 295, 309, 319, 329, 330, 251, 287, 375, 36]. ending [375]. England [95, 90, 155, 325, 385, 41, 188, 218, 115, 391, 98, 232, 107, 312, 242]. English [94, 121, 169, 181, 236, 278, 293, 315, 358, 283, 391]. Enteric [337]. Entitled [216]. Entrance [197, 251, 287, 330, 375]. Especial [321]. Esq [385, 214, 261, 48, 297, 342]. Establishing [91]. Europe [134, 209, 272, 302]. European [133, 98, 371]. Evils [108]. Excess [306]. Exchange [131]. Exchanges [3]. Executions [218]. Exeter [7]. Exhibiting [312]. Exhibition [191, 276]. Expenditure [155, 91]. Expense [202]. Exports [356, 190, 273]. Extended [178]. Extension [175]. Extent [326]. Extract [89]. Extracts [299].

F [300, 105, 299, 201, 344, 15, 47, 147, 304, 387]. F.R.S. [214, 261, 48]. Factory [265]. Failures [325]. Fall [353]. Families [79]. Famine [120]. Famines [333, 354]. Farmer [372]. Feasibility [366]. Fellows [274]. Fermented [53]. Fever [337]. Fifth [376, 375, 201]. Finance [101, 153, 4]. Financial [10, 63, 100, 97, 127, 143, 193, 197, 244, 251, 284, 287, 292, 323, 330, 355, 367, 375, 36, 289]. Fire [235, 277, 314, 357, 289]. Fires [182, 235, 277, 314, 357, 289]. First [198, 188, 200, 352, 47, 197]. Food [326]. Foreign [393, 190, 273, 326, 322, 203, 321, 80, 3, 371]. Foresters [274]. Form [99]. Fortieth [19, 144, 143, 15]. Forty [198, 252, 251, 288, 287, 331, 330, 376, 375, 263, 105, 299, 201, 344, 47, 387, 197, 147]. Forty-Eighth [344]. Forty-Fifth [376, 375, 201]. Forty-First [198, 47, 197].

Forty-Fourt [330]. Forty-Fourth [331, 147]. Forty-Ninth [387]. Forty-Second [252, 251]. Forty-Seventh [299]. Forty-Sixth [263]. Forty-Third [288, 287, 105]. Fourt [330]. Fourth [331, 352, 147]. France [150, 170, 185, 225, 393, 370, 98, 312]. Free [16, 4]. Friendly [189, 274]. furnished [87]. Future [79, 176].

G. [297, 342]. Gardens [205]. Gas [243]. Gazette [380]. General [100, 127, 193, 244, 284, 282, 323, 367, 178, 255]. Generations [79]. Geographical [41, 215]. Geography [352]. German [312, 139, 194]. Germany [128, 185]. Given [203]. Giving [91]. Glasgow [163, 263]. Going [203]. Goods [393]. Government [192, 99, 303, 283, 145, 240, 254]. Governmental [91]. Governments [248]. Gradutes [224]. Grain [206]. Great [166, 16, 313, 35, 80]. Growing [273].

Hamburg [80]. Health [191, 52, 175]. Held [10, 7, 63, 97, 143, 191, 197, 251, 276, 330, 73, 263, 105, 201, 344, 264, 262, 352, 36, 147, 346, 19, 287, 375, 163, 231, 299, 298, 387]. Helena [65]. Her [125]. High [108, 40]. Higher [107]. Historical [289]. History [100, 127, 193, 244, 284, 323, 355, 367, 216]. Home [125, 11]. Hon [140]. Hospitals [282]. House [68, 197, 251, 287, 375, 160, 130, 131, 245, 368, 391]. Howard [216, 78, 216]. Hungarian [266].

II [354]. Imports [306, 190, 273]. Imposed [174]. Improved [205, 175, 74]. Inaugural [385, 45, 72, 104, 160, 214, 297]. Incidence [12]. Inclusive [222, 203]. Income [124]. Increase [133, 391]. Increasing [253]. India [183, 247, 221]. Indian [101]. Indians [25]. Influence [393, 17]. Information [1, 22, 30, 44, 51, 71, 77, 103, 113, 159, 173, 213, 229, 260, 271, 296, 310, 341, 351, 384, 46].

Inland [80]. Inquiry [107]. Institutions [189, 230, 24, 242]. Insurance [129, 289]. Intellectual [41]. Interest [239, 379, 107, 40]. Interference [53]. International [119, 191, 276, 335, 303, 241, 352, 3, 73, 262, 298, 352]. Intoxicating [174]. Investigated [230]. Ireland [366, 391, 35, 80]. Isles [377]. Issue [74, 312]. Issued [203]. Italian [153, 303]. Italy [312]. Items [224]. IV [231, 163].

J [297, 342]. James

[10, 63, 97, 143, 45, 72, 104, 36, 214, 261]. January [194]. John [78, 216]. Joint [180, 316, 2]. July [224, 352]. June [10, 63, 197, 36, 10, 63, 97, 142, 143, 196, 250, 251, 286, 287, 329, 330, 375, 36]. Jupiter [360]. Justice [115].

King [197, 251, 287, 330, 375, 160].

Kingdom [135, 165, 207, 282, 316, 381, 66, 221, 273, 230, 311, 321].

Lady [94]. Lady-Day [94]. Land

[110, 205, 146, 98, 232]. Landowner [372]. Languages [377]. Larger [209]. Last [370, 300, 218]. Law [132, 156, 80]. Laws [239, 31, 232]. Leading [392]. Learned [68]. Led [107]. Leeds [48]. Lefevre [297, 342]. Legal [230, 115]. Legislation [220, 92]. Legislative [53]. Letters [89]. Libraries [16]. Library [123, 142, 158, 172, 187, 196, 212, 228, 238, 250, 259, 270, 281, 286, 295, 309, 294, 319, 329, 339, 340, 350, 396, 383, 362, 374, 395, 109]. Limited [392]. Limits [53]. Liquors [174, 53]. List [203]. Literature [121, 181, 236, 278, 315, 358]. Little [98]. Liverpool [19, 15, 264]. Lloyd [168, 184, 234, 256, 285, 324, 369, 65, 203]. Loan [99]. Local [95, 164, 283, 145, 46, 35, 12, 39]. London [197, 235, 251, 277, 287, 314, 330, 357, 375, 45, 72, 104, 160]. Loss [65, 203]. Losses [168, 129, 184, 234, 256, 285, 372, 65]. Lower [54].

M [214]. M.A [261]. M.P. [385, 297, 342]. Majesty [125]. Manchester [334]. Manufactures [183]. Manufacturing [313]. March [187, 238, 281, 319]. Margin [392]. Marine [168, 129, 184, 234, 256, 285, 324, 369]. Market [93]. Marriages [188, 200, 302]. Massachusetts [224]. Mathematical [162, 327]. may [392, 313]. Mayr [276]. Means [99]. Mechanic [24]. Medical [352]. Meeting [19, 10, 7, 63, 97, 144, 143, 198, 197, 252, 251, 288, 287, 331, 330, 376, 375, 263, 105, 299, 201, 344, 15, 47, 36, 387, 197, 147]. Melada [222]. Members [224]. Memorandum [221]. Mental [167]. Mercantile [138, 177, 203]. Merchant [65]. Metropolis [182, 243, 391]. Metropolitan [235, 277, 314, 357]. Migrations [199]. Military [52]. Milk [337]. Mineral [135, 165]. Miscellanea [8, 13, 20, 28, 33, 37, 42, 49, 56, 61, 69, 75]. Miscellaneous [381]. Molasses [222]. Monetary [11, 40]. Money [93]. Monsieur [118]. Morals [175]. Mortality [282]. Motions [360]. Movement [134, 146, 371]. Movements [393]. Mr [140]. Mr. [81, 89, 301]. Much [98]. Municipal [240]. Murder [218].

National [114, 91, 264, 312, 289, 163, 231, 346]. Nations [145, 203]. Nature [326]. Naval [52]. Navies [138, 177]. Navigation [59]. Netherlands [312]. Neutralising [108]. Nineteenth [231]. Ninth [7, 97, 262, 298, 387, 276]. Non [230]. Non-Legal-Tender [230]. North [25]. Northern [334]. Norwich [106]. Note [200, 230, 312]. Notes [294, 350, 349, 383, 362, 361, 373, 382, 374, 395, 16, 85, 80].

Notices [308, 338, 318]. November

[386, 385, 214, 261, 297, 45, 72, 104, 160, 342]. Number [203, 391].

Obstacles [136]. Occupied [205, 98].

October [48, 163, 231, 106, 264, 346]. Odd [274]. Office [111]. Official [59]. Old [139]. Opening [163, 231, 261, 106, 15, 47, 342]. Operation [74, 254]. Operations [178]. Operative [146, 136]. Orchards [205]. Orders [274]. Other [253, 321, 98]. Our [190, 11, 326, 91]. Owens [334]. Owing [322]. Owned [205]. Owners [98].

Palgrave [81]. Paper [81, 300, 216]. Paris [348, 336, 240, 352, 215]. Parliament [388, 32]. Parliamentary [91]. Part [354]. Past [282, 190, 333, 354]. Pasture [205]. Pauper [94]. Pauperism [202, 31]. Peerage [248]. Peerages [248]. Penal [220]. Percentage [203]. Periodical [9, 14, 21, 29, 34, 38, 43, 50, 57, 62, 70, 76, 84, 96, 102, 112, 122, 141, 157, 171, 186, 195, 211, 227, 237, 249, 258, 269, 280, 320, 328, 363]. Periods [174]. Permanent [119, 335, 352]. Pesth [276, 262, 298]. Petersburg [73]. Phases [365]. Physical [167, 265]. Place [218, 216]. Planet [360]. Playfair [255]. Plymouth [300, 299]. Political [162, 327]. Poor [132, 156]. Population [154, 167, 209, 370, 300, 79, 302]. Populations [290]. Portion [120]. Ports [317]. Position [150, 364]. Positions [312]. Post [111]. Posted [65, 203]. Powers [88]. Preliminary [262]. Preparation [32]. Preponderance [273]. Present [224, 282, 364, 333, 354]. Presented [10, 63, 97, 143, 197, 251, 287, 330, 375, 36]. President [385, 163, 263, 231, 105, 299, 201, 214, 261, 106, 344, 15, 264, 47, 48, 147, 346, 297, 342, 387]. Price [393, 17, 108]. Prices [128, 380, 356, 353]. Principal [109, 302]. Principles [239, 220, 219]. Printing [32]. Prison [54, 241, 352]. Procedure [115].

Proceedings [5, 18, 10, 63, 97, 144, 143, 198, 197, 252, 251, 288, 287, 331, 330, 376, 386, 375, 36]. Produced [321]. Producing [206]. Production [136]. Profits [316, 253]. Progress [101, 226, 114, 190, 275, 364, 302, 321]. Promotion [163, 231, 264, 346]. Property [91, 124]. Proposed [304]. Prospective [253]. Protectionist [321]. Province [205]. Public [191, 23]. Published [303]. Purchase [88, 85].

Qualities [41]. Quarter [123, 142, 158, 172, 187, 196, 212, 228, 238, 250, 259, 270, 281, 286, 295, 309, 319, 329, 396]. Quasi [322]. Quasi-Sovereign [322]. Question [365, 98]. Quetelet [118].

Races [359]. Railroads [207, 208]. Railway [87]. Railways [88, 90, 148, 169, 128, 293, 221, 86, 253, 85]. Rate [179, 360, 239, 107, 40]. Rates [149, 174, 161, 107, 199]. Read [300]. Reason [313]. Receipts [155]. Recognition [23]. Recommendations [6]. Reconstruction [124]. Reference [23, 321]. References [312]. Registered [203]. Registrar [178]. Registrar-General [178]. Registration [140]. Regulate [239]. Reichsbank [312]. Relating [224, 65, 391, 242, 217]. Relation [107]. Relative [185, 79]. Relief [202]. Religious [242]. Remarks [343, 232, 80, 312]. Report [10, 63, 97, 151, 143, 140, 197, 251, 287, 276, 330, 375, 303, 73, 262, 298, 36]. Reports [255, 52, 352]. Requirements [265]. Reserve [107]. Reserves [230]. Resource [305]. Respect [6]. Respectively [203]. Results [100, 127, 193, 244, 284, 323, 367, 370, 26, 272]. Returns [9, 14, 21, 29, 34, 38, 43, 50, 57, 62, 70, 76, 83, 84, 96, 102, 112, 166, 122, 141, 157, 171, 186, 195, 211, 227, 237, 249, 258, 269, 307, 280, 347, 320, 328, 389, 368, 363, 391]. Review [100]. Right [140]. Rise [128]. Roll [94]. Rooms [10, 63, 97, 143, 197, 251, 287, 330, 375, 45, 72, 104, 160, 36]. Russia [290, 312, 371]. Russian [292, 67].

Safely [392]. Safety [191]. Sailing [203]. Sale [53]. Sanitary [161]. Savings [303]. Science [291, 263, 105, 299, 23, 201, 106, 15, 47, 48, 147, 304, 387, 19, 7, 163, 263, 231, 105, 299, 201, 15, 264, 47, 147, 346]. Sciences [23]. Sclater [140]. Sclater-Booth [140]. Scotland [4, 391]. Sea [203]. Sea-Going [203]. Seamen [52]. Second [252, 251, 298, 352, 346]. Section [263, 105, 299, 201, 344, 15, 264, 47, 147, 387, 300, 304]. Security [99]. Separately [91]. September [19, 262, 158, 212, 259, 295, 276, 263, 105, 15, 298]. Serving [125]. Session [5, 352]. Sessional [10, 63, 97, 143, 197, 251, 287, 330, 375, 36]. Seventh [299, 36]. Seventy [218]. Several [248]. shall [91]. Shaw [297, 342]. Sheffield [387]. Showing [203]. Shown [52, 91]. Silver [268, 365]. Similar [189]. since [321]. Sixth [10, 263]. So-Called [23]. Social [163, 231, 23, 264, 48, 346, 106]. Societies [68, 189, 274]. Society [5, 18, 10, 6, 63, 109, 97, 143, 197, 251, 287, 330, 375, 385, 45, 72, 104, 160, 214, 261, 36, 297, 342]. Soldiers [52]. Some [209, 225, 224, 395, 365, 300, 115, 65, 302, 379, 274, 98, 232, 107, 80]. Somerset [197, 251, 287, 375, 160]. Sound [91]. Sources [326]. South [192, 226]. Sovereign [322]. Spain [223]. Special [282, 23]. Spot [359]. Spots [279]. Square [10, 63, 97, 143, 45, 72, 104, 36]. St. [10, 63, 97, 143, 73, 45, 72, 104, 65, 36]. State [88, 85, 312]. Statement [205]. States [206, 224, 322, 272, 83, 208, 222, 385, 176]. Statist [78]. Statist. [216]. Statistical [10, 119, 152, 197, 224, 276, 349, 335, 336, 361, 373, 382, 370, 16, 52, 385, 73, 263, 46, 214, 261, 26, 343, 272, 262, 298, 352, 377, 74, 297, 342, 289, 5, 18, 6, 63, 97, 143, 251, 287, 330, 375, 36]. Statistically [230]. Statistics [64, 82, 168, 135, 165, 120, 184, 222, 225, 267, 266, 234, 256, 285, 292, 324, 369, 381, 394, 388, 345, 59, 303, 116, 105, 299, 60, 201, 92, 115, 65, 15, 67, 125, 54, 241, 47, 274, 147, 32, 215, 304, 387, 242, 254, 217, 87, 219]. Statists [192]. Steamers [203]. Steps [248]. Stock [180, 316, 2, 131]. Stockholm [352]. Strand [197, 251, 287, 330, 375]. Strength [185]. Subsequently [203]. Suez [137, 65]. Sugar [222]. Suggested [360]. Suggestions [99, 91, 175, 46, 108, 32, 74]. Suicide [125]. Summary [67]. Sun [359, 279]. Sun-Spot [359]. Supplement [216]. Supplies [326, 79]. Supply [337, 243, 27]. Support [23, 242]. Supremacy [313]. Survey [377]. Sweden [149, 80]. Swiss [111]. System [91, 74, 312].

Table [248, 203]. Tabular [378]. Taken [218]. Tariffs [321]. Tax [110, 124]. Taxation [95, 164, 35, 12, 39]. Telegraphs [253]. Telegraphy [64]. Ten [17, 125]. Tender [230]. Termed [99]. Terminating [253]. Terrace [197, 251, 287, 375, 160]. Test [161]. Their [188, 175, 31, 289]. Them [174]. Theory [162, 327]. there [313]. thereon [312]. think [313]. Third [288, 287, 105]. Thirty [10, 7, 63, 97, 36]. Thirty-Eighth [63]. Thirty-Ninth [7, 97]. Thirty-Seventh [36]. Thirty-Sixth [10]. Thomas [385]. Thursday [10, 63, 36]. Times [282, 89]. Tobacco [66]. Tonnage [60]. Total [203]. Touching [65]. Town [16, 79]. Town-Libraries [16]. Towns [209]. Trade [151, 204, 183, 317, 393, 59, 190, 273, 163, 231, 264, 48, 321, 346, 332]. Traffic [137, 221, 65]. Travellers [371]. Treaties [272]. Treaty [317]. Troops [125]. True [216]. Tuesday [385, 45, 72, 104, 160, 214, 261, 297, 342]. Turkey [290]. Turkish [305]. Twentieth [264]. Twenty [190, 391, 346]. Twenty-Second [346]. Tyler [87].

Uncalled [392]. Underwriter [82].

Underwriting [257]. United

[224, 83, 135, 165, 207, 208, 222, 282, 316, 381, 66, 221, 273, 385, 230, 311, 321, 176]. Universal [203]. University [359, 334]. Unlimited [392]. Upon [321, 255, 174, 17, 203, 32, 199]. Use [343].

Validity [254]. Valuation [86, 253]. Value [91, 161]. Variations [3]. Various [174, 203]. Veritas [203]. Vessels [65, 203]. Vict [88]. Victoria [364, 217]. Violent [90]. Vital [219]. Volume [1, 22, 30, 44, 51, 71, 77, 103, 113, 159, 173, 213, 229, 260, 271, 296, 310, 341, 351, 384].

W.C [197, 251, 287]. Wages [134, 128, 170]. Wales [325, 218, 242]. was [105, 201, 147]. Water [253]. Wheat [210]. whence [326]. Which [147, 239, 91, 105, 201, 313, 107]. William [48]. Wool [27]. Workings [221]. Works [327, 349, 361, 373, 382]. World [138, 177, 203, 333, 354]. Wrecks [279].

Year [10, 63, 97, 143, 182, 197, 235, 256, 251, 257, 277, 285, 287, 314, 324, 330, 357, 369, 375, 221, 36]. Years [129, 190, 17, 353, 218, 177, 203, 125, 391, 107, 3].

Zealand [275].

References

Anonymous:1870:VIa

[1] Anonymous. Volume information. Journal of the Statistical Society of London, 33(1):i–vii, March 1870. CODEN ???? ISSN 0959-5341. URL http://

                                             www.jstor.org/stable/2338864.

Levi:1870:JSC

[2] Leone Levi. On joint stock companies. Journal of the Statistical Society of London, 33(1):1–41, March 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338865.

Seyd:1870:ICV

[3] Ernest Seyd. On international coinage and the variations of the foreign exchanges during recent years. Journal of the Statistical Society of London, 33(1): 42–73, March 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338866.

Buchanan:1870:FFC

[4] Robert Buchanan. Finance of the free Church of Scotland. Journal of the Statistical Society of London, 33(1):74–110, March 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338867.

Anonymous:1870:PSSa

[5] Anonymous. Proceedings of the Statistical Society, session 1868–69. Journal of the Statistical Society of London, 33(1):111–112, March 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338868.

Anonymous:1870:RCS

[6] Anonymous. Recommendations of the Council of the Statistical Society with respect to the census of 1871. Journal of the Statistical Society of London, 33(1): 113, March 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338869.

Anonymous:1870:TNM

[7] Anonymous. Thirty-ninth meeting of the British Association for the Advancement of Science, held at exeter, 18th– 25th August, 1869. Journal of the Statistical Society of London, 33(1):114–115, March 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338870.

Anonymous:1870:Ma

[8] Anonymous. Miscellanea. Journal of the Statistical Society of London, 33(1):116– 169, March 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338871.

Anonymous:1870:PRa

[9] Anonymous. Periodical returns. Journal of the Statistical Society of London, 33(1):170–203, March 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338872.

Anonymous:1870:RCF

[10] Anonymous. Report of the Council for the financial year ended 31st December, 1869, and for the sessional year ended with June, 1870, presented at the thirtysixth anniversary meeting of The Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Thursday, 23rd June, 1870; with the proceedings of that meeting. Journal of the Statistical Society of London, 33(2):205–215, June 1870. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338714.

Patterson:1870:OHM

[11] R. H. Patterson. On our home monetary drains, and the crisis of 1866. Journal of the Statistical Society of London, 33(2): 216–242, June 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338715.

Rogers:1870:ILT

[12] James E. Thorold Rogers. On the incidence of local taxation. Journal of the Statistical Society of London, 33(2):243– 263, June 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338716.

Anonymous:1870:Mb

[13] Anonymous. Miscellanea. Journal of the Statistical Society of London, 33(2):264– 287, June 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338717.

Anonymous:1870:PRb

[14] Anonymous. Periodical returns. Journal of the Statistical Society of London, 33(2):288–308, June 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338718.

Jevons:1870:OAP

[15] W. Stanley Jevons. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the fortieth meeting, at Liverpool, September, 1870. Journal of the Statistical Society of London, 33(3): 309–326, September 1870. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338902.

Axon:1870:SNF

[16] William E. A. Axon. Statistical notes on the free town-libraries of Great Britain and the continent. Journal of the Statistical Society of London, 33(3):327–365, September 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338903.

Forwood:1870:IPU

[17] William B. Forwood. The influence of price upon the cultivation and consumption of cotton during the ten years 1860–70. Journal of the Statistical Society of London, 33(3):366–383, September 1870. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338904.

Anonymous:1870:PSSb

[18] Anonymous. Proceedings of the Statistical Society. Journal of the Statistical Society of London, 33(3):384–385, September 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338905.

Anonymous:1870:FMB

[19] Anonymous. Fortieth meeting of the British Association for the Advancement of Science, held at Liverpool, 14th–21st September, 1870. Journal of the Statistical Society of London, 33(3):386–387, September 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338906.

Anonymous:1870:Mc

[20] Anonymous. Miscellanea. Journal of the Statistical Society of London, 33(3): 388–411, September 1870. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338907.

Anonymous:1870:PRc

[21] Anonymous. Periodical returns. Journal of the Statistical Society of London, 33 (3):412–431, September 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338908.

Anonymous:1870:VIb

[22] Anonymous. Volume information. Journal of the Statistical Society of London, 33(4):553–568, December 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338845.

Guy:1870:CSP

[23] William A. Guy. On the claims of science to public recognition and support; with special reference to the so-called “social sciences.”. Journal of the Statistical Society of London, 33(4):433–451, December 1870. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338846.

Renals:1870:MIE

[24] E. Renals. On mechanics’ institutions and the elementary education bill. Journal of the Statistical Society of London, 33(4):452–455, December 1870. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338847.

Heywood:1870:ANA

[25] James Heywood. On the aptitude of North American Indians for agriculture. Journal of the Statistical Society of London, 33(4):456–462, December 1870. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338848.

Hill:1870:SRC

[26] Berkeley Hill. Statistical results of the Contagious Diseases Acts. Journal of the Statistical Society of London, 33(4): 463–485, December 1870. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338849.

Hamilton:1870:WS

[27] Archibald Hamilton. On wool supply. Journal of the Statistical Society of London, 33(4):486–521, December 1870. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338850.

Anonymous:1870:Md

[28] Anonymous. Miscellanea. Journal of the Statistical Society of London, 33(4): 522–532, December 1870. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338851.

Anonymous:1870:PRd

[29] Anonymous. Periodical returns. Journal of the Statistical Society of London, 33 (4):533–552, December 1870. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338852.

Anonymous:1871:VIa

[30] Anonymous. Volume information. Journal of the Statistical Society of London, 34(1):i–vii, March 1871. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338757.

Seyd:1871:CLT

[31] Ernest Seyd. On currency laws and their effects on pauperism. Journal of the Statistical Society of London, 34(1):1– 20, March 1871. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338758.

Purdy:1871:SUP

[32] Frederick Purdy. Suggestions upon the preparation and printing of Parliament statistics. Journal of the Statistical Society of London, 34(1):21–56, March 1871. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338759.

Anonymous:1871:Ma

[33] Anonymous. Miscellanea. Journal of the Statistical Society of London, 34(1):57– 82, March 1871. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338760.

Anonymous:1871:PRa

[34] Anonymous. Periodical returns. Journal of the Statistical Society of London, 34(1):83–109, March 1871. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338761.

Palgrave:1871:LTG

[35] R. H. Inglis Palgrave. On the local taxation of Great Britain and Ireland. Journal of the Statistical Society of London, 34(2):111–235, June 1871. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338656.

Newmarch:1871:RCF

[36] William Newmarch. Report of the Council for the financial year ended 31st December, 1870, an for the sessional year ended with June, 1871, presented at the thirty-seventh anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Thursday, 22nd June, 1871; with the proceedings of that meeting. Journal of the Statistical Society of London, 34(2): 236–251, June 1871. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338657.

Anonymous:1871:Mb

[37] Anonymous. Miscellanea. Journal of the Statistical Society of London, 34(2):252– 260, June 1871. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338658.

Anonymous:1871:PRb

[38] Anonymous. Periodical returns. Journal of the Statistical Society of London, 34(2):261–280, June 1871. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338659.

Scott:1871:LT

[39] John Scott. On local taxation. Journal of the Statistical Society of London, 34 (3):281–333, September 1871. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338825.

Patterson:1871:RIE

[40] R. H. Patterson. On the rate of interest,-and the effects of a high bankrate during commercial and monetary crises. Journal of the Statistical Society of London, 34(3):334–356, September 1871. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338826.

Clarke:1871:GDI

[41] Hyde Clarke. On the geographical distribution of intellectual qualities in England. Journal of the Statistical Society of London, 34(3):357–373, September 1871. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338827.

Anonymous:1871:Mc

[42] Anonymous. Miscellanea. Journal of the Statistical Society of London, 34(3): 374–383, September 1871. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338828.

Anonymous:1871:PRc

[43] Anonymous. Periodical returns. Journal of the Statistical Society of London, 34 (3):384–407, September 1871. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338829.

Anonymous:1871:VIb

[44] Anonymous. Volume information. Journal of the Statistical Society of London, 34(4):539–553, December 1871. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338784.

Farr:1871:IAD

[45] William Farr. Inaugural address delivered at the Society’s rooms, 12, St. James’s Square, London, on Tuesday, 21st November, 1871. Journal of the Statistical Society of London, 34(4):409– 423, December 1871. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338785.

Hammick:1871:SCA

[46] James T. Hammick. Suggestions for the collection and arrangement of local statistical information. Journal of the Statistical Society of London, 34(4): 424–460, December 1871. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338786.

Neaves:1871:OAP

[47] Lord Neaves. Opening address of the President of Section F (Economic Science and Statistics), of the British Association for the Advancement of Science, at the forty-first meeting, at Edinburgh, August, 1871. Journal of the Statistical Society of London, 34(4):461–475, December 1871. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338787.

Newmarch:1871:AWN

[48] William Newmarch. Address by William Newmarch, Esq., F.R.S., as President of the Economy and Trade Department, Social Science Association, at Leeds, 10th October, 1871. Journal of the Statistical Society of London, 34(4):476– 498, December 1871. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338788.

Anonymous:1871:Md

[49] Anonymous. Miscellanea. Journal of the Statistical Society of London, 34(4): 499–511, December 1871. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338789.

Anonymous:1871:PRd

[50] Anonymous. Periodical returns. Journal of the Statistical Society of London, 34 (4):512–538, December 1871. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338790.

Anonymous:1872:VIa

[51] Anonymous. Volume information. Journal of the Statistical Society of London, 35(1):i–vii, March 1872. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338687.

Balfour:1872:CHS

[52] T. Graham Balfour. Comparative health of seamen and soldiers, as shown by the Naval and military statistical reports. Journal of the Statistical Society of London, 35(1):1–24, March 1872. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338688.

Levi:1872:LLI

[53] Leone Levi. On the limits of legislative interference with the sale of fermented liquors. Journal of the Statistical Society of London, 35(1):25–56, March 1872. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338689.

Mouat:1872:PDS

[54] Frederic John Mouat. On prison discipline and statistics in lower Bengal. Journal of the Statistical Society of London, 35(1):57–106, March 1872. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338690.

Hamilton:1872:C

[55] Archibald Hamilton. On the colonies. Journal of the Statistical Society of London, 35(1):107–126, March 1872. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338691.

Anonymous:1872:Ma

[56] Anonymous. Miscellanea. Journal of the Statistical Society of London, 35(1):127– 150, March 1872. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338692.

Anonymous:1872:PRa

[57] Anonymous. Periodical returns. Journal of the Statistical Society of London, 35(1):151–170, March 1872. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338693.

Chubb:1872:BAC

[58] Hammond Chubb. The Bank Act and the crisis of 1866. Journal of the Statistical Society of London, 35(2):171–195, June 1872. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338745.

Bourne:1872:OTN

[59] Stephen Bourne. The official trade and navigation statistics. Journal of the Statistical Society of London, 35(2):196– 217, June 1872. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338746.

Glover:1872:TSD

[60] John Glover. Tonnage statistics of the decade 1860–70. Journal of the Statistical Society of London, 35(2):218–230, June 1872. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338747.

Anonymous:1872:Mb

[61] Anonymous. Miscellanea. Journal of the Statistical Society of London, 35(2):231– 244, June 1872. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338748.

Anonymous:1872:PRb

[62] Anonymous. Periodical returns. Journal of the Statistical Society of London, 35(2):245–264, June 1872. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338749.

Anonymous:1872:RCF

[63] Anonymous. Report of the Council for the financial year ended 31st December, 1871, and for the sessional year ended with June, 1872, presented at the thirtyeighth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on Thursday, 20th June, 1872; with the proceedings of that meeting. Journal of the Statistical Society of London, 35(3):265–271, September 1872. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338679.

Anderson:1872:ST

[64] James Anderson. Statistics of telegraphy. Journal of the Statistical Society of London, 35(3):272–326, September 1872. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338680.

Jeula:1872:SSR

[65] Henry Jeula. Some statistics relating to the traffic through the Suez Canal; to merchant vessels touching at St. Helena; and to losses posted on “Lloyd’s Loss Book”. Journal of the Statistical Society of London, 35(3):327–333, September 1872. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338681.

Axon:1872:CTU

[66] William E. A. Axon. On the consumption of tobacco in the United Kingdom, 1801–70. Journal of the Statistical Society of London, 35(3):334–340, September 1872. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338682.

Michell:1872:SSR

[67] Robert Michell. Summary of statistics of the Russian empire. Journal of the Statistical Society of London, 35 (3):341–372, September 1872. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338683.

Anonymous:1872:HAL

[68] Anonymous. House accomodation for learned societies. Journal of the Statistical Society of London, 35(3):373–375, September 1872. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338684.

Anonymous:1872:Mc

[69] Anonymous. Miscellanea. Journal of the Statistical Society of London, 35(3): 376–388, September 1872. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338685.

Anonymous:1872:PRc

[70] Anonymous. Periodical returns. Journal of the Statistical Society of London, 35 (3):389–416, September 1872. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338686.

Anonymous:1872:VIb

[71] Anonymous. Volume information. Journal of the Statistical Society of London, 35(4):565–582, December 1872. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338762.

Farr:1872:IAD

[72] William Farr. Inaugural address delivered at the Society’s rooms, 12, St. James’s Square, London, on Tuesday, 19th November, 1872. Journal of the Statistical Society of London, 35(4):417– 430, December 1872. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338763.

Brown:1872:REI

[73] Samuel Brown. Report on the eighth International Statistical Congress, held at St. Petersburg, 22nd/10th August to 29th/17th, 1872. Journal of the Statistical Society of London, 35(4):431–457, December 1872. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338764.

Seyd:1872:SCO

[74] Ernest Seyd. Statistical critique on the operation of the Bank Charter Act of 1844, and suggestions for an improved system of issue. Journal of the Statistical Society of London, 35(4):458–540, December 1872. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338765.

Anonymous:1872:Md

[75] Anonymous. Miscellanea. Journal of the Statistical Society of London, 35(4): 541–545, December 1872. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338766.

Anonymous:1872:PRd

[76] Anonymous. Periodical returns. Journal of the Statistical Society of London, 35 (4):546–564, December 1872. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338767.

Anonymous:1873:VIa

[77] Anonymous. Volume information. Journal of the Statistical Society, 36(1):i– vii, March 1873. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338802.

Guy:1873:JHS

[78] William A. Guy. John Howard as statist. Journal of the Statistical Society, 36(1): 1–18, March 1873. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338803. See supplement

[216].

Galton:1873:RST

[79] Francis Galton. The relative supplies from town and country families, to the population of future generations. Journal of the Statistical Society, 36(1):19– 26, March 1873. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338804.

Palgrave:1873:NBG

[80] Robert Harry Inglis Palgrave. Notes on banking in Great Britain and Ireland, Sweden, Denmark, and Hamburg; with some remarks on the amount of bills in circulation, both inland and foreign, in Great Britain and Ireland; and the Banking Law of Sweden. Journal of the Statistical Society, 36(1):27–152, March 1873. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/

2338805. See discussion [81].

Anonymous:1873:DMP

[81] Anonymous. Discussion on Mr. Palgrave’s paper. Journal of the Statistical Society, 36(1):153–157, March 1873. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/

2338806. See [80].

Anonymous:1873:US

[82] Anonymous. Underwriter’s statistics. Journal of the Statistical Society, 36(1): 157–158, March 1873. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338807.

Anonymous:1873:ARU

[83] Anonymous. Agricultural returns of the United States. Journal of the Statistical Society, 36(1):158, March 1873. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338808.

Anonymous:1873:PRa

[84] Anonymous. Periodical returns. Journal of the Statistical Society, 36(1):159– 176, March 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338809.

Martin:1873:NPR

[85] R. Biddulph Martin. Notes on the purchase of the railways by the state. Journal of the Statistical Society, 36(2):177– 255, June 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338909.

Farr:1873:VR

[86] W. Farr. On the valuation of railways. Journal of the Statistical Society, 36(2): 256–259, June 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338910.

Tyler:1873:RSF

[87] Captain Tyler. Railway statistics furnished by Captain Tyler. Journal of the Statistical Society, 36(2):259–270, June 1873. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338911.

Anonymous:1873:PSP

[88] Anonymous. Powers of state to purchase railways. clauses of 7 and 8 vict., cap. 85 (9th August, 1844). Journal of the Statistical Society, 36(2):270–272, June 1873. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338912.

Anonymous:1873:EMC

[89] Anonymous. Extract from Mr. Chubb’s letters to the “Times”. Journal of the Statistical Society, 36(2):272–274, June 1873. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338913.

Anonymous:1873:VDE

[90] Anonymous. Violent deaths in England connected with railways, 1871. Journal of the Statistical Society, 36(2):275– 276, June 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338914.

Fellows:1873:ONP

[91] Frank P. Fellows. On our national parliamentary accounts, with suggestions for establishing a Doomsday book giving the value of national governmental property or assets as the basis of a sound system of accounts, by which expenditure for capital and current account shall be separately shown. Journal of the Statistical Society, 36(2):277– 302, June 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338915.

Janson:1873:SL

[92] Frederick H. Janson. The statistics of legislation. Journal of the Statistical Society, 36(2):303–320, June 1873. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338916.

Anonymous:1873:DMM

[93] Anonymous. The dangers of the money market. Journal of the Statistical Society, 36(2):321–324, June 1873. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338917.

Anonymous:1873:EPR

[94] Anonymous. The English pauper roll at Lady-day. Journal of the Statistical Society, 36(2):324–326, June 1873. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338918.

Anonymous:1873:LTE

[95] Anonymous. The local taxation of England. Journal of the Statistical Society, 36(2):326, June 1873. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338919.

Anonymous:1873:PRb

[96] Anonymous. Periodical returns. Journal of the Statistical Society, 36(2):327– 346, June 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338920.

Anonymous:1873:RCF

[97] Anonymous. Report of the Council for the financial year ended 31st December, 1872, and for the sessional year ended with June, 1873, presented at the thirtyninth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on 30th June, 1873; with the proceedings of that meeting. Journal of the Statistical Society, 36 (3):347–361, September 1873. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338708.

Norman:1873:AAQ

[98] George Warde Norman. An attempt to answer the question, why it is that so much land is occupied and cultivated by its owners in France, and some other European countries on the continent, and so little in England? Journal of the Statistical Society, 36(3):362–374, September 1873. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338709.

Barron:1873:SGL

[99] Henry Barron. Suggestions for a government loan by means of a new form of security to be termed “accumulative bonds”. Journal of the Statistical Society, 36(3):375–385, September 1873. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338710.

Anonymous:1873:GRC

[100] Anonymous. General results of the commercial and financial history and review of 1872. Journal of the Statistical Society, 36(3):386–432, September 1873. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338711.

Anonymous:1873:PIF

[101] Anonymous. The progress of Indian finance. Journal of the Statistical Society, 36(3):433–438, September 1873. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338712.

Anonymous:1873:PRc

[102] Anonymous. Periodical returns. Journal of the Statistical Society, 36(3):439–465, September 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338713.

Anonymous:1873:VIb

[103] Anonymous. Volume information. Journal of the Statistical Society, 36(4):607– 629, December 1873. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338669.

Guy:1873:IAD

[104] William A. Guy. Inaugural address delivered at the Society’s rooms, 12, St. James’s Square, London, on Tuesday, 18th November, 1873. Journal of the Statistical Society, 36(4):467–485, December 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338670.

Forster:1873:APS

[105] W. E. Forster. Address of the President of Section F (Economic Science and Statistics) of the British Association for the Advancement of Science, at the forty-third meeting, which was held at Bradford, September, 1873. Journal of the Statistical Society, 36(4):486–501, December 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338671.

Houghton:1873:OAP

[106] Lord Houghton. Opening address of the President of the Social Science Congress, at Norwich, October, 1873. Journal of the Statistical Society, 36(4):502–528, December 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338672.

Palgrave:1873:RBR

[107] R. H. Inglis Palgrave. On the relation of the banking reserve of the Bank of England to the current rate of interest, with an inquiry into some of the causes which have led to the higher rates charged in recent years. Journal of the Statistical Society, 36(4):529–564, December 1873. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338673.

Hill:1873:HPC

[108] Rowland Hill. High price of coal. suggestions for neutralising its evils. Journal of the Statistical Society, 36(4):565–579, December 1873. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338674.

Anonymous:1873:PAL

[109] Anonymous. Principal additions to the Library of the Society during 1873. Journal of the Statistical Society, 36(4):580– 581, December 1873. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338675.

Anonymous:1873:LT

[110] Anonymous. The land tax. Journal of the Statistical Society, 36(4):581–583, December 1873. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338676.

Anonymous:1873:SPO

[111] Anonymous. The Swiss post Office. Journal of the Statistical Society, 36(4): 583–584, December 1873. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338677.

Anonymous:1873:PRd

[112] Anonymous. Periodical returns. Journal of the Statistical Society, 36(4):585–605, December 1873. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338678.

Anonymous:1874:VIa

[113] Anonymous. Volume information. Journal of the Statistical Society, 37(1):i– viii, March 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338830.

Baxter:1874:RPN

[114] R. Dudley Baxter. The recent progress of national debts. Journal of the Statistical Society, 37(1):1–20, March 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338831.

Janson:1874:SSC

[115] F. H. Janson. Some statistics of the courts of justice and of legal procedure in England. Journal of the Statistical Society, 37(1):21–47, March 1874. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338832.

Cork:1874:SAB

[116] Nathaniel Cork. Statistics of Australasian banking. Journal of the Statistical Society, 37(1):48–68, March 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338833.

Beverley:1874:CB

[117] Henry Beverley. The census of Bengal. Journal of the Statistical Society, 37(1):69–113, March 1874. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338834.

Anonymous:1874:MQ

[118] Anonymous. Monsieur Quetelet. Journal of the Statistical Society, 37(1):114– 115, March 1874. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338835.

Anonymous:1874:PCI

[119] Anonymous. The Permanent Commission of the International Statistical Congress. Journal of the Statistical Society, 37(1):115–117, March 1874. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338836.

Anonymous:1874:SPA

[120] Anonymous. Statistics of a portion of the area of the famine in Bengal. Journal of the Statistical Society, 37(1):117– 123, March 1874. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338837.

Anonymous:1874:EL

[121] Anonymous. English literature of 1873. Journal of the Statistical Society, 37(1): 123–125, March 1874. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338838.

Anonymous:1874:PRa

[122] Anonymous. Periodical returns. Journal of the Statistical Society, 37(1):126– 150, March 1874. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338839.

Anonymous:1874:ALDa

[123] Anonymous. Additions to the library during the quarter. Journal of the Statistical Society, 37(1):151–153, March 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338840.

Levi:1874:RIP

[124] Leone Levi. On the reconstruction of the income and property tax. Journal of the Statistical Society, 37(2):155– 186, June 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338719.

Millar:1874:SDS

[125] W. H. Millar. Statistics of deaths by suicide among Her Majesty’s British troops serving at home and abroad during the ten years 1862–71. Journal of the Statistical Society, 37(2):187–192, June 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338720.

Martin:1874:E

[126] John Biddulph Martin. The elections of 1868 and 1874. Journal of the Statistical Society, 37(2):193–230, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338721.

Anonymous:1874:GRC

[127] Anonymous. General results of the commercial and financial history of 1873. Journal of the Statistical Society, 37(2): 230–241, June 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338722.

Anonymous:1874:RGR

[128] Anonymous. Railways in Germany and the rise of wages and prices, 1867–73. Journal of the Statistical Society, 37(2): 241–242, June 1874. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338723.

Anonymous:1874:MIL

[129] Anonymous. Marine insurance losses in recent years. Journal of the Statistical Society, 37(2):242–244, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338724.

Anonymous:1874:BCH

[130] Anonymous. The Banker’s Clearing House. Journal of the Statistical Society, 37(2):245, June 1874. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338725.

Anonymous:1874:SEC

[131] Anonymous. The Stock Exchange Clearing House. Journal of the Statistical Society, 37(2):245–251, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338726.

Anonymous:1874:PLAa

[132] Anonymous. Poor law administration in 1871–72. Journal of the Statistical Society, 37(2):251–255, June 1874. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338727.

Anonymous:1874:IEA

[133] Anonymous. The increase of European armaments. Journal of the Statistical Society, 37(2):255–257, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338728.

Anonymous:1874:MAW

[134] Anonymous. The movement of agricultural wages in Europe. Journal of the Statistical Society, 37(2):257–258, June 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338729.

Anonymous:1874:MSUa

[135] Anonymous. Mineral statistics of the United Kingdom. Journal of the Statistical Society, 37(2):258–260, June 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338730.

Anonymous:1874:OCO

[136] Anonymous. Obstacles to co-operative production. Journal of the Statistical Society, 37(2):260–264, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338731.

Anonymous:1874:TSC

[137] Anonymous. The traffic of the Suez Canal. Journal of the Statistical Society, 37(2):264, June 1874. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338732.

Anonymous:1874:MNW

[138] Anonymous. The mercantile Navies of the world. Journal of the Statistical Society, 37(2):264–265, June 1874. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338733.

Anonymous:1874:NOG

[139] Anonymous. The new and old German coinages. Journal of the Statistical Society, 37(2):266–270, June 1874. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338734.

Anonymous:1874:RDR

[140] Anonymous. Report of deputation to the Right Hon. Mr. Sclater-Booth, on the Registration of Births and Deaths Bill.

Journal of the Statistical Society, 37(2): 270–273, June 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338735.

Anonymous:1874:PRb

[141] Anonymous. Periodical returns. Journal of the Statistical Society, 37(2):274– 293, June 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338736.

Anonymous:1874:ALDb

[142] Anonymous. Additions to the library during the quarter ended 30th June, 1874. Journal of the Statistical Society, 37(2):294–298, June 1874. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338737.

Anonymous:1874:RCF

[143] Anonymous. Report of the Council for the financial year ended 31st December, 1873, and for the sessional year ended 30th June, 1874, presented at the fortieth anniversary meeting of the Statistical Society, held at the Society’s rooms, 12, St. James’s Square, on the 30th of June, 1874, with the proceedings of that meeting. Journal of the Statistical Society, 37(3):299–309, September 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338921.

Anonymous:1874:PFA

[144] Anonymous. Proceedings of the fortieth anniversary meeting. Journal of the Statistical Society, 37(3):310–312, September 1874. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338922.

Dilke:1874:LGA

[145] Charles W. Dilke. Local government among different nations. Journal of the Statistical Society, 37(3):313–326, September 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338923.

Brabrook:1874:COL

[146] E. W. Brabrook. The co-operative land movement. Journal of the Statistical Society, 37(3):327–337, September 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338924.

OHagan:1874:APS

[147] Lord O’Hagan. Address of the President of Section F (Economic Science and Statistics) of the British Association for the Advancement of Science, at the Forty-Fourth Meeting, which was held at Belfast, August, 1874. Journal of the Statistical Society, 37(3):338–349, September 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338925.

Anonymous:1874:AR

[148] Anonymous. American railways. Journal of the Statistical Society, 37(3):350– 355, September 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338926.

Anonymous:1874:DRS

[149] Anonymous. Death-rates in Sweden.

Journal of the Statistical Society, 37(3): 355–356, September 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338927.

Anonymous:1874:CPF

[150] Anonymous. The commercial position of France. Journal of the Statistical Society, 37(3):356–359, September 1874. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338928.

Anonymous:1874:CRT

[151] Anonymous. The customs report on the trade of 1873. Journal of the Statistical Society, 37(3):359–365, September 1874. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338929.

Anonymous:1874:SA

[152] Anonymous. Statistical abstracts. Journal of the Statistical Society, 37(3):365– 370, September 1874. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338930.

Anonymous:1874:IF

[153] Anonymous. Italian finance. Journal of the Statistical Society, 37(3):370–374, September 1874. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338931.

Anonymous:1874:EP

[154] Anonymous. The effective population. Journal of the Statistical Society, 37(3): 374–376, September 1874. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338932.

Anonymous:1874:CRE

[155] Anonymous. County receipts and expenditure in England. Journal of the Statistical Society, 37(3):376–377, September 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338933.

Anonymous:1874:PLAb

[156] Anonymous. Poor law administration in 1873. Journal of the Statistical Society, 37(3):377–384, September 1874. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338934.

Anonymous:1874:PRc

[157] Anonymous. Periodical returns. Journal of the Statistical Society, 37(3):385–405, September 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338935.

Anonymous:1874:ALDc

[158] Anonymous. Additions to the library during the quarter ended 30th September, 1874. Journal of the Statistical Society, 37(3):406–410, September 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338936.

Anonymous:1874:VIb

[159] Anonymous. Volume information. Journal of the Statistical Society, 37(4):571– 595, December 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338694.

Guy:1874:IAD

[160] William A. Guy. Inaugural address delivered at the Society’s rooms, Somerset House Terrace, King’s College, London, on Tuesday, 17th November, 1874. Journal of the Statistical Society, 37(4): 411–436, December 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338695.

Humphreys:1874:VDR

[161] N. A. Humphreys. The value of deathrates as a test of sanitary condition. Journal of the Statistical Society, 37(4): 437–477, December 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338696.

Anonymous:1874:MTP

[162] Anonymous. The mathematical theory of political economy. Journal of the Statistical Society, 37(4):478–488, December 1874. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338697.

Campbell:1874:OAP

[163] George Campbell. Opening address of the President of the Department. IV, “Economy and trade,” of the National Association for the Promotion of Social Science, at the eighteenth annual Congress, held at Glasgow, October, 1874. Journal of the Statistical Society, 37(4):489–505, December 1874. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338698.

Anonymous:1874:LT

[164] Anonymous. Local taxation. Journal of the Statistical Society, 37(4):506–510, December 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338699.

Anonymous:1874:MSUb

[165] Anonymous. Mineral statistics of the United Kingdom. Journal of the Statistical Society, 37(4):510–516, December 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338700.

Anonymous:1874:ARG

[166] Anonymous. Agricultural returns of Great Britain. Journal of the Statistical Society, 37(4):516–518, December 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338701.

Anonymous:1874:PMC

[167] Anonymous. Physical and mental condition of the convict population. Journal of the Statistical Society, 37(4):518–519, December 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338702.

Anonymous:1874:LSM

[168] Anonymous. Lloyd’s statistics of marine losses in 1872. Journal of the Statistical Society, 37(4):519–531, December 1874. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338703.

Anonymous:1874:ER

[169] Anonymous. English railways in 1873.

Journal of the Statistical Society, 37(4): 532–540, December 1874. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338704.

Anonymous:1874:WF

[170] Anonymous. Wages in France. Journal of the Statistical Society, 37(4):540–542, December 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338705.

Anonymous:1874:PRd

[171] Anonymous. Periodical returns. Journal of the Statistical Society, 37(4):543–562, December 1874. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338706.

Anonymous:1874:ALDd

[172] Anonymous. Additions to the library during the quarter ended 31st December, 1874. Journal of the Statistical Society, 37(4):563–569, December 1874. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338707.

Anonymous:1875:VIa

[173] Anonymous. Volume information. Journal of the Statistical Society, 38(1):i– viii, March 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338887.

Burns:1875:CIL

[174] Dawson Burns. The consumption of intoxicating liquors at various periods as affected by the rates of duty imposed upon them. Journal of the Statistical Society, 38(1):1–32, March 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338888.

Gatliff:1875:IDT

[175] Charles Gatliff. On improved dwellings and their beneficial effect on health and morals, with suggestions for their extension. Journal of the Statistical Society, 38(1):33–63, March 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338889.

Norman:1875:FUS

[176] George Warde Norman. The future of the United States. Journal of the Statistical Society, 38(1):64–78, March 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338890.

Jeula:1875:MNW

[177] Henry Jeula. The mercantile Navies of the world in the years 1870 and 1874 compared. Journal of the Statistical Society, 38(1):79–86, March 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338891.

Anonymous:1875:EOR

[178] Anonymous. Extended operations of the Registrar-General. Journal of the Statistical Society, 38(1):87–88, March 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338892.

Anonymous:1875:DR

[179] Anonymous. The death-rate. Journal of the Statistical Society, 38(1):88– 89, March 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338893.

Anonymous:1875:JSB

[180] Anonymous. Joint stock banks. Journal of the Statistical Society, 38(1):89– 90, March 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338894.

Anonymous:1875:EL

[181] Anonymous. English literature in 1874.

Journal of the Statistical Society, 38(1): 90–92, March 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338895.

Anonymous:1875:FMD

[182] Anonymous. Fires in the Metropolis during the year 1874. Journal of the Statistical Society, 38(1):92–96, March 1875. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338896.

Anonymous:1875:TMI

[183] Anonymous. The trade and manufactures of India. Journal of the Statistical Society, 38(1):96–98, March 1875. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338897.

Anonymous:1875:LSM

[184] Anonymous. Lloyd’s statistics of marine losses in 1873. Journal of the Statistical Society, 38(1):98–111, March 1875. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338898.

Anonymous:1875:RSF

[185] Anonymous. The relative strength of France and Germany. Journal of the Statistical Society, 38(1):112–118, March 1875. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338899.

Anonymous:1875:PRa

[186] Anonymous. Periodical returns. Journal of the Statistical Society, 38(1):119– 145, March 1875. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338900.

Anonymous:1875:ALDa

[187] Anonymous. Additions to the library during the quarter ended 31st March, 1875. Journal of the Statistical Society, 38(1):146–151, March 1875. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338901.

Darwin:1875:MBF

[188] George H. Darwin. Marriages between first cousins in England and their effects. Journal of the Statistical Society, 38(2):153–184, June 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338660.

Brabrook:1875:FSS

[189] E. W. Brabrook. Friendly societies and similar institutions. Journal of the Statistical Society, 38(2):185–214, June 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338661.

Bourne:1875:POF

[190] Stephen Bourne. The progress of our foreign trade, imports and exports, during the past twenty years. Journal of the Statistical Society, 38(2):215–248, June 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338662.

Anonymous:1875:IEC

[191] Anonymous. International exhibition and Congress of public health and safety, to be held in Brussels in 1876. Journal of the Statistical Society, 38(2):249– 252, June 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338663.

Anonymous:1875:CGS

[192] Anonymous. Conference of government statists, South Australia. Journal of the Statistical Society, 38(2):252–257, June 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338664.

Anonymous:1875:GRC

[193] Anonymous. General results of the commercial and financial history of 1874. Journal of the Statistical Society, 38(2): 257–266, June 1875. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338665.

Anonymous:1875:GBA

[194] Anonymous. German Bank Act of the 30th January, 1875. Journal of the Statistical Society, 38(2):267–282, June 1875. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338666.

Anonymous:1875:PRb

[195] Anonymous. Periodical returns. Journal of the Statistical Society, 38(2):283– 302, June 1875. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338667.

Anonymous:1875:ALDb

[196] Anonymous. Additions to the library during the quarter ended 30th June, 1875. Journal of the Statistical Society, 38(2):303–308, June 1875. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338668.

Anonymous:1875:RCF

[197] Anonymous. Report of the Council for the financial year ended 31st December, 1874, and for the sessional year ended 29th June, 1875, presented at the Forty-First Anniversary Meeting of The Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College Entrance), Strand, W.C., London, on the 29th of June, 1875, with the proceedings of that meeting. Journal of the Statistical Society, 38(3):309–317, September 1875. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338768.

Anonymous:1875:PFF

[198] Anonymous. Proceedings of the fortyfirst anniversary meeting. Journal of the Statistical Society, 38(3):318–323, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338769.

Welton:1875:EMU

[199] Thomas A. Welton. The effect of migrations upon death-rates. Journal of the Statistical Society, 38(3):324–343, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338770.

Darwin:1875:NMF

[200] George H. Darwin. Note on the marriages of first cousins. Journal of the Statistical Society, 38(3):344–348, September 1875. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2338771.

Heywood:1875:APS

[201] James Heywood. Address of the President of Section F (Economic Science and Statistics) of the British Association for the Advancement of Science, at the forty-fifth meeting, which was held at Bristol, August, 1875. Journal of the Statistical Society, 38(3):349–360, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338772.

Anonymous:1875:PER

[202] Anonymous. Pauperism and the expense of relief. Journal of the Statistical Society, 38(3):361–363, September 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338773.

Jeula:1875:TSN

[203] Henry Jeula. A table showing the number of vessels, British and foreign, posted on “Lloyd’s Loss Book,” for eighteen years, from 1857 to 1874 inclusive; with the percentage upon total British registered vessels from 1857 to 1869 inclusive; and, subsequently the percentage of sailing vessels and steamers respectively so posted, upon the mercantile sea-going sailing vessels and steamers of various nations of the world as given in the “Universal Mercantile List” issued by the Bureau Veritas. Journal of the Statistical Society, 38(3):364–367, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338774.

Anonymous:1875:TCB

[204] Anonymous. Trade and condition of Belgium. Journal of the Statistical Society, 38(3):367–369, September 1875. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338775.

Anonymous:1875:SAL

[205] Anonymous. Statement of the area of land owned, area occupied, improved and under crops, pasture, gardens, and orchards, in each province of Canada, according to the census of 1870–71. Journal of the Statistical Society, 38(3): 370, September 1875. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338776.

Anonymous:1875:GPS

[206] Anonymous. Grain producing states. Journal of the Statistical Society, 38(3): 371, September 1875. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338777.

Anonymous:1875:RUK

[207] Anonymous. The railroads of the United Kingdom. Journal of the Statistical Society, 38(3):371–375, September 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338778.

Anonymous:1875:RUS

[208] Anonymous. The railroads of the United States. Journal of the Statistical Society, 38(3):375–379, September 1875. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338779.

Anonymous:1875:PSL

[209] Anonymous. Population of some of the larger towns of Europe. Journal of the Statistical Society, 38(3):379–382, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338780.

Anonymous:1875:WC

[210] Anonymous. The wheat crop of 1875.

Journal of the Statistical Society, 38(3): 382–384, September 1875. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338781.

Anonymous:1875:PRc

[211] Anonymous. Periodical returns. Journal of the Statistical Society, 38(3):385–405, September 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338782.

Anonymous:1875:ALDc

[212] Anonymous. Additions to the library during the quarter ended 30th September, 1875. Journal of the Statistical Society, 38(3):406–411, September 1875. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338783.

Anonymous:1875:VIb

[213] Anonymous. Volume information. Journal of the Statistical Society, 38(4):569– 591, December 1875. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338959.

Heywood:1875:IAJ

[214] James Heywood. The inaugural address of James Heywood, Esq., M. A., F.R.S., President of the Statistical Society, delivered in Tuesday, the 16th November, 1875. Journal of the Statistical Society, 38(4):413–421, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338960.

Ravenstein:1875:SPG

[215] E. G. Ravenstein. Statistics at the Paris Geographical Congress. Journal of the Statistical Society, 38(4):422–429, December 1875. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338961.

Guy:1875:JHT

[216] William A. Guy. John Howard’s true place in history. a supplement to the paper entitled “John Howard as Statist.”. Journal of the Statistical Society, 38(4): 430–437, December 1875. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338962. See [78].

Stark:1875:SRC

[217] William Emery Stark. Statistics relating to the Colony of Victoria, Australia. Journal of the Statistical Society, 38(4):438–462, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338963.

Guy:1875:EMT

[218] William A. Guy. On the executions for murder that have taken place in England and Wales during the last seventy years. Journal of the Statistical Society, 38(4):463–486, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338964.

Verwey:1875:PVS

[219] A. J. Verwey. Principles of vital statistics. Journal of the Statistical Society, 38 (4):487–513, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338965.

Burt:1875:PPL

[220] John T. Burt. Principles of penal legislation. Journal of the Statistical Society, 38(4):514–516, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338966.

Baynes:1875:MCT

[221] J. A. Baynes. Memorandum on the cost, traffic, and workings of the railways in the United Kingdom and in India, for the year 1874, compared with 1873. Journal of the Statistical Society, 38(4):516–526, December 1875. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338967.

Anonymous:1875:SCS

[222] Anonymous. Statistics of the consumption of sugar, melada, and molasses in the United States from 1860 to 1873 inclusive. Journal of the Statistical Society, 38(4):526–532, December 1875. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338968.

Anonymous:1875:CS

[223] Anonymous. The commerce of Spain. Journal of the Statistical Society, 38(4): 532–536, December 1875. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338969.

Anonymous:1875:SSI

[224] Anonymous. Some statistical items relating to the gradutes and present members of Amherst College, Massachusetts, United States of America, July, 1875. Journal of the Statistical Society, 38(4): 536–537, December 1875. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338970.

Anonymous:1875:SES

[225] Anonymous. Some early statistics of France. Journal of the Statistical Society, 38(4):537–538, December 1875. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338971.

Anonymous:1875:PSA

[226] Anonymous. Progress of South Australia. Journal of the Statistical Society, 38(4):538–539, December 1875. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338972.

Anonymous:1875:PRd

[227] Anonymous. Periodical returns. Journal of the Statistical Society, 38(4):540–559, December 1875. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338973.

Anonymous:1875:ALDd

[228] Anonymous. Additions to the library during the quarter ended 31st December, 1875. Journal of the Statistical Society, 38(4):560–568, December 1875. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338974.

Anonymous:1876:VIa

[229] Anonymous. Volume information. Journal of the Statistical Society, 39(1):i– viii, March 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339039.

Dun:1876:BIB

[230] John Dun. The banking institutions, bullion reserves, and non-legal-tender note circulation of the United Kingdom statistically investigated. Journal of the Statistical Society, 39(1):1–197, March 1876. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339040.

Duff:1876:OAP

[231] M. E. Grant Duff. Opening address of the President of the Department IV, “Economy and Trade,” of the National Association for the Promotion of Social Science, at the nineteenth annual Congress, held at Brighton, in October, 1875. Journal of the Statistical Society, 39(1):198–210, March 1876. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339041.

Norman:1876:SRL

[232] George Warde Norman. Some remarks on the Land Laws of England. Journal of the Statistical Society, 39(1):211– 218, March 1876. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339042.

Anonymous:1876:CAC

[233] Anonymous. The climate, agriculture, commerce, &c., of Egypt. Journal of the Statistical Society, 39(1):218–223, March 1876. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339043.

Anonymous:1876:LSMa

[234] Anonymous. Lloyd’s statistics of marine losses in 1874. Journal of the Statistical Society, 39(1):223–237, March 1876. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339044.

Anonymous:1876:FLD

[235] Anonymous. The fires in London during the year 1875, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 39(1):238–241, March 1876. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339045.

Anonymous:1876:EL

[236] Anonymous. English literature in 1875. Journal of the Statistical Society, 39(1): 241–243, March 1876. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339046.

Anonymous:1876:PRa

[237] Anonymous. Periodical returns. Journal of the Statistical Society, 39(1):244– 270, March 1876. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339047.

Anonymous:1876:ALDa

[238] Anonymous. Additions to the library during the quarter ended 31st March, 1876. Journal of the Statistical Society, 39(1):271–276, March 1876. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339048.

Baxter:1876:PWR

[239] Robert Baxter. On the principles which regulate the rate of interest and on the currency laws. Journal of the Statistical Society, 39(2):277–298, June 1876. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339116.

Dilke:1876:MGP

[240] Charles W. Dilke. On the municipal government of Paris. Journal of the Statistical Society, 39(2):299–310, June 1876. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339117.

Mouat:1876:IPS

[241] Frederic J. Mouat. On international prison statistics. Journal of the Statistical Society, 39(2):311–331, June 1876. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339118.

Skeats:1876:SRS

[242] Herbert S. Skeats. Statistics relating to the support of religious institutions in England and Wales. Journal of the Statistical Society, 39(2):332–349, June 1876. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339119.

Chubb:1876:SGM

[243] Harry Chubb. The supply of gas to the Metropolis. Journal of the Statistical Society, 39(2):350–380, June 1876. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339120.

Anonymous:1876:GRC

[244] Anonymous. General results of the commercial and financial history of 1875. Journal of the Statistical Society, 39(2): 381–391, June 1876. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339121.

Anonymous:1876:BCH

[245] Anonymous. The Bankers’ Clearing House. Journal of the Statistical Society, 39(2):392, June 1876. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339122.

Anonymous:1876:DB

[246] Anonymous. The Domesday Book of 1873. Journal of the Statistical Society, 39(2):393–410, June 1876. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339123.

Anonymous:1876:CBI

[247] Anonymous. The census of British India of 1871–72. Journal of the Statistical Society, 39(2):411–416, June 1876. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339124.

Anonymous:1876:TPS

[248] Anonymous. A table of Peerages, steps in the Peerage, and Baronetcies, conferred by the several governments from 1830–74. Journal of the Statistical Society, 39(2):417, June 1876. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339125.

Anonymous:1876:PRb

[249] Anonymous. Periodical returns. Journal of the Statistical Society, 39(2):418– 437, June 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339126.

Anonymous:1876:ALDb

[250] Anonymous. Additions to the library during the quarter ended 30th June, 1876. Journal of the Statistical Society, 39(2):438–443, June 1876. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339127.

Anonymous:1876:RCF

[251] Anonymous. Report of the Council for the financial year ended 31st December, 1875, and for the sessional year ended 27th June, 1876, presented at the forty-second anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, W.C., London, on the 27th of June, 1876, with the proceedings of that meeting. Journal of the Statistical Society, 39(3):445–458, September 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338937.

Anonymous:1876:PFS

[252] Anonymous. Proceedings of the fortysecond anniversary meeting. Journal of the Statistical Society, 39(3):459–463, September 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338938.

Farr:1876:VRT

[253] William Farr. On the valuation of railways, telegraphs, water companies, canals, and other commercial concerns, with prospective, deferred, increasing, decreasing, or terminating profits. Journal of the Statistical Society, 39(3):464– 539, September 1876. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338939.

Stansfeld:1876:VAG

[254] James Stansfeld. On the validity of the annual government statistics of the operation of the Contagious Diseases Acts. Journal of the Statistical Society, 39(3): 540–572, September 1876. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338940.

Anonymous:1876:CGP

[255] Anonymous. Consul-General Playfair’s reports upon Algeria. Journal of the Statistical Society, 39(3):573–577, September 1876. CODEN ???? ISSN 0959-

                                        5341. URL http://www.jstor.org/
                                        stable/2338941.

Anonymous:1876:LSMb

[256] Anonymous. Lloyd’s statistics of marine losses for the year 1875. Journal of the Statistical Society, 39(3):577–591, September 1876. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338942.

Anonymous:1876:UY

[257] Anonymous. Underwriting in the year 1805. Journal of the Statistical Society, 39(3):592–593, September 1876. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338943.

Anonymous:1876:PRc

[258] Anonymous. Periodical returns. Journal of the Statistical Society, 39(3):594–614, September 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338944.

Anonymous:1876:ALDc

[259] Anonymous. Additions to the library during the quarter ended 30th September, 1876. Journal of the Statistical Society, 39(3):615–619, September 1876. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338945.

Anonymous:1876:VIb

[260] Anonymous. Volume information. Journal of the Statistical Society, 39(4):771– 794, December 1876. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339154.

Heywood:1876:OAJ

[261] James Heywood. The opening address of James Heywood, Esq., M.A., F.R.S., President of the Statistical Society, delivered on Tuesday, the 21st November, 1876. Journal of the Statistical Society, 39(4):621–627, December 1876. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339155.

Mouat:1876:PRN

[262] Frederic J. Mouat. Preliminary report of the ninth International Statistical Congress, held at Buda-Pesth, from 1st to 7th September, 1876. Journal of the Statistical Society, 39(4):628–647, December 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339156.

Campbell:1876:APE

[263] George Campbell. Address of the President of the economic science and statistical section of the British Association for the Advancement of Science, at the forty-sixth meeting, held at Glasgow, September, 1876. Journal of the Statistical Society, 39(4):648–660, December 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339157.

Lefevre:1876:APD

[264] G. J. Shaw Lefevre. Address of the President of the Department of “Economy and Trade” Section National Association for the Promotion of Social Science, at the Twentieth annual Congress, held at Liverpool, in October, 1876. Journal of the Statistical Society, 39(4):661–680, December 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339158.

Roberts:1876:PRF

[265] Charles Roberts. The physical requirements of factory children. Journal of the Statistical Society, 39(4):681–733, December 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339159.

Anonymous:1876:HS

[266] Anonymous. Hungarian statistics. Journal of the Statistical Society, 39(4):734– 739, December 1876. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339160.

Anonymous:1876:CS

[267] Anonymous. Condensed statistics. Journal of the Statistical Society, 39(4):739– 741, December 1876. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2339161.

Anonymous:1876:DS

[268] Anonymous. The depreciation of silver. Journal of the Statistical Society, 39 (4):741–744, December 1876. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339162.

Anonymous:1876:PRd

[269] Anonymous. Periodical returns. Journal of the Statistical Society, 39(4):745–764, December 1876. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339163.

Anonymous:1876:ALDd

[270] Anonymous. Additions to the library during the quarter ended 31st December, 1876. Journal of the Statistical Society, 39(4):765–770, December 1876. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339164.

Anonymous:1877:VIa

[271] Anonymous. Volume information. Journal of the Statistical Society, 40(1):i– viii, March 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339212.

Levi:1877:SRR

[272] Leone Levi. Statistical results of the recent treaties of commerce between different states of Europe. Journal of the Statistical Society, 40(1):1–18, March 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339213.

Bourne:1877:GPI

[273] Stephen Bourne. The growing preponderance of imports over exports in the foreign and colonial trade of the United Kingdom. Journal of the Statistical Society, 40(1):19–41, March 1877. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339214.

Neison:1877:SSA

[274] Francis G. P. Neison. Some statistics of the affiliated orders of Friendly Societies (Odd Fellows and Foresters). Journal of the Statistical Society, 40(1):42– 89, March 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339215.

Hamilton:1877:REP

[275] Archibald Hamilton. On the recent economic progress of New Zealand. Journal of the Statistical Society, 40(1):90– 130, March 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339216.

Anonymous:1877:RDM

[276] Anonymous. Report by Dr. Mayr on the exhibition of charts and diagrams at the Ninth International Statistical Congress, held at Buda-Pesth, in September, 1876. Journal of the Statistical Society, 40(1): 131–136, March 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339217.

Anonymous:1877:FLD

[277] Anonymous. The fires of London during the year 1876, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 40(1):136–140, March 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339218.

Anonymous:1877:EL

[278] Anonymous. English literature in 1876.

Journal of the Statistical Society, 40(1): 140–142, March 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339219.

Anonymous:1877:WSS

[279] Anonymous. Wrecks and Sun spots.

Journal of the Statistical Society, 40(1): 143–144, March 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339220.

Anonymous:1877:PR

[280] Anonymous. Periodical returns. Journal of the Statistical Society, 40(1):145– 170, March 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339221.

Anonymous:1877:ALDa

[281] Anonymous. Additions to the library during the quarter ended 31st March, 1877. Journal of the Statistical Society, 40(1):171–176, March 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339222.

Anonymous:1877:MHG

[282] Anonymous. The mortality of hospitals, general and special, in the United Kingdom, in times past and present. Journal of the Statistical Society, 40(2):177– 261, June 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339196.

Craigie:1877:CEL

[283] P. G. Craigie. The cost of English local government. Journal of the Statistical Society, 40(2):262–296, June 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339197.

Anonymous:1877:GRC

[284] Anonymous. General results of the commercial and financial history of 1876. Journal of the Statistical Society, 40(2): 297–308, June 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339198.

Anonymous:1877:LSM

[285] Anonymous. Lloyd’s statistics of marine losses for the year 1876. Journal of the Statistical Society, 40(2):308–323, June 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339199.

Anonymous:1877:ALDb

[286] Anonymous. Additions to the library during the quarter ended 30th June, 1877. Journal of the Statistical Society, 40(2):324–331, June 1877. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339200.

Anonymous:1877:RCF

[287] Anonymous. Report of the Council for the financial year ended 31st December, 1876, and for the sessional year ended 26th June, 1877, presented at the fortythird anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, W.C., London, on the 26th of June, 1877, with the proceedings of that meeting. Journal of the Statistical Society, 40(3):333–341, September 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339072.

Anonymous:1877:PFT

[288] Anonymous. Proceedings of the fortythird anniversary meeting. Journal of the Statistical Society, 40(3):342–346, September 1877. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339073.

Walford:1877:FFI

[289] Cornelius Walford. Fires and fire insurance considered under their historical, financial, statistical, and national aspects. Journal of the Statistical Society, 40(3):347–432, September 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339074.

Ravenstein:1877:PRT

[290] E. G. Ravenstein. The populations of Russia and Turkey. Journal of the Statistical Society, 40(3):433–467, September 1877. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339075.

Anonymous:1877:ESB

[291] Anonymous. Economic science and the British Association. Journal of the Statistical Society, 40(3):468–476, September 1877. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339076.

Anonymous:1877:RFS

[292] Anonymous. Russian financial statistics. Journal of the Statistical Society, 40 (3):476–485, September 1877. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339077.

Anonymous:1877:ER

[293] Anonymous. English railways in 1876.

Journal of the Statistical Society, 40(3):

485–497, September 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339078.

Anonymous:1877:NRA

[294] Anonymous. Notes on recent additions to the library. Journal of the Statistical Society, 40(3):497–500, September 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339079.

Anonymous:1877:ALDc

[295] Anonymous. Additions to the library during the quarter ended 30th September, 1877. Journal of the Statistical Society, 40(3):501–508, September 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339080.

Anonymous:1877:VIb

[296] Anonymous. Volume information. Journal of the Statistical Society, 40(4):679– 705, December 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339010.

Shaw-Lefevre:1877:IAG

[297] G. J. Shaw-Lefevre. Inaugural address of G. J. Shaw-Lefevre, Esq., M.P., President of the Statistical Society, delivered on Tuesday, the 20th November, 1877. Journal of the Statistical Society, 40(4):509–530, December 1877. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339011.

Mouat:1877:SCR

[298] Frederic J. Mouat. Second and concluding report of the ninth International Statistical Congress, held in Buda-Pesth, in September, 1876. Journal of the Statistical Society, 40(4):531–556, December 1877. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339012.

Fortescue:1877:EAP

[299] Earl Fortescue. Extracts from the address of the President of Section F, “Economic Science and Statistics,” of the British Association for the Advancement of Science, at the forty-seventh meeting, held at Plymouth, August, 1877. Journal of the Statistical Society, 40(4): 557–567, December 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339013.

Farr:1877:SDP

[300] William Farr. On some doctrines of population. A paper read in Section F of the British Association, at Plymouth, in August last. Journal of the Statistical Society, 40(4):568–581, December 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339014.

Anonymous:1877:MCA

[301] Anonymous. Mr. Caird’s address at Aberdeen. Journal of the Statistical Society, 40(4):582–592, December 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339015.

Martin:1877:BDM

[302] Frederick Martin. Births, deaths, and marriages, and the comparative progress of population in some of the principal countries of Europe. Journal of the Statistical Society, 40(4):593–613, December 1877. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339016.

Brabrook:1877:ARP

[303] Edward W. Brabrook. Abstract of a report published by the Italian government on the international statistics of savings banks. Journal of the Statistical Society, 40(4):614–630, December 1877. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339017.

S:1877:PDS

[304] F. R. S. The proposed discontinuance of Section F, Economic Science and Statistics, at the British Association. Journal of the Statistical Society, 40(4):631–633, December 1877. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339018.

Anonymous:1877:TR

[305] Anonymous. Turkish resource. Journal of the Statistical Society, 40(4):633–645, December 1877. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339019.

Anonymous:1877:EI

[306] Anonymous. The excess of imports. Journal of the Statistical Society, 40(4): 645–650, December 1877. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339020.

Anonymous:1877:AR

[307] Anonymous. Agricultural returns for 1877. Journal of the Statistical Society, 40(4):651–659, December 1877. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339021.

Anonymous:1877:NNB

[308] Anonymous. Notices of new books. Journal of the Statistical Society, 40(4):

660–667, December 1877. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339022.

Anonymous:1877:ALDd

[309] Anonymous. Additions to the library during the quarter ended 31st December, 1877. Journal of the Statistical Society, 40(4):668–677, December 1877. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339023.

Anonymous:1878:VIa

[310] Anonymous. Volume information. Journal of the Statistical Society, 41(1):i– viii, March 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339201.

Giffen:1878:RAC

[311] Robert Giffen. Recent accumulations of capital in the United Kingdom. Journal of the Statistical Society, 41(1):1– 39, March 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339202.

Seyd:1878:DEP

[312] Ernest Seyd. Diagrams exhibiting the positions of The Bank of England, The Bank of France, The German Reichsbank, The National Bank of Austria, The Netherlands Bank, The National Bank of Belgium, The National Bank of Italy, and The State Bank of Russia; with remarks thereon and references to the note issue system of The Bank of England. Journal of the Statistical Society, 41(1):40–86, March 1878. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339203.

Mundella:1878:WCW

[313] A. J. Mundella. What are the conditions on which the commercial and manufacturing supremacy of Great Britain depend, and is there any reason to think they have been, or may be, endangered? Journal of the Statistical Society, 41(1):87–134, March 1878. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339204.

Anonymous:1878:FLD

[314] Anonymous. The fires of London during the year 1877, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 41(1):135–138, March 1878. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339205.

Anonymous:1878:EL

[315] Anonymous. English literature in 1877. Journal of the Statistical Society, 41(1): 138–141, March 1878. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339206.

Anonymous:1878:PJS

[316] Anonymous. The profits of joint stock banking in the United Kingdom. Journal of the Statistical Society, 41(1):141– 148, March 1878. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339207.

Anonymous:1878:TCT

[317] Anonymous. Trade of the Chinese treaty ports. Journal of the Statistical Society, 41(1):149–150, March 1878. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339208.

Block:1878:NNB

[318] Maurice Block. Notices of new books.

Journal of the Statistical Society, 41(1): 151–160, March 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339209.

Anonymous:1878:ALDa

[319] Anonymous. Additions to the library during the quarter ended 31st March, 1878. Journal of the Statistical Society, 41(1):161–166, March 1878. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339210.

Anonymous:1878:PRa

[320] Anonymous. Periodical returns. Journal of the Statistical Society, 41(1):167– 186, March 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339211.

Newmarch:1878:PFT

[321] William Newmarch. On the progress of the foreign trade of the United Kingdom since 1856, with especial reference to the effects produced upon it by the protectionist tariffs of other countries. Journal of the Statistical Society, 41(2):187– 298, June 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339094.

Clarke:1878:DSQ

[322] Hyde Clarke. On the debts of sovereign and quasi-sovereign states, owing by foreign countries. Journal of the Statistical Society, 41(2):299–347, June 1878. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339095.

Anonymous:1878:GRC

[323] Anonymous. General results of the commercial and financial history of 1877. Journal of the Statistical Society, 41(2): 348–366, June 1878. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339096.

Anonymous:1878:LSM

[324] Anonymous. Lloyd’s statistics of marine casualties for the year 1877. Journal of the Statistical Society, 41(2):366– 383, June 1878. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339097.

Anonymous:1878:FEW

[325] Anonymous. Failures in England and Wales. Journal of the Statistical Society, 41(2):384–386, June 1878. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339098.

Bourne:1878:NEO

[326] Stephen Bourne. The nature and extent of our foreign food supplies, and the sources from whence they are derived. Journal of the Statistical Society, 41(2):386–398, June 1878. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339099.

Anonymous:1878:BWM

[327] Anonymous. Bibliography of works on the mathematical theory of political economy. Journal of the Statistical Society, 41(2):398–401, June 1878. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339100.

Anonymous:1878:PRb

[328] Anonymous. Periodical returns. Journal of the Statistical Society, 41(2):401– 407, June 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339101.

Anonymous:1878:ALDb

[329] Anonymous. Additions to the library during the quarter ended 30th June, 1878. Journal of the Statistical Society, 41(2):408–414, June 1878. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339102.

Anonymous:1878:RCF

[330] Anonymous. Report of the Council for the financial year ended 31st December, 1877, and for the sessional year ended 25th June, 1878, presented at the fortyfourt anniversary meeting of the Statistical Society, held at the Society’s rooms, King’s College entrance, Strand, London, on the 25th of June, 1878, and of the proceedings at the meeting. Journal of the Statistical Society, 41(3):415–424, September 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338975.

Anonymous:1878:PFF

[331] Anonymous. Proceedings of the fortyfourth anniversary meeting. Journal of the Statistical Society, 41(3):425–426, September 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338976.

Shaw-Lefevre:1878:DT

[332] G. J. Shaw-Lefevre. The depression of trade. Journal of the Statistical Society, 41(3):427–432, September 1878. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338977.

Walford:1878:FWP

[333] Cornelius Walford. The famines of the world: Past and present. Journal of the Statistical Society, 41(3):433–535, September 1878. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338978.

Heywood:1878:OCM

[334] James Heywood. The Owens College, Manchester, and a Northern University. Journal of the Statistical Society, 41(3): 536–548, September 1878. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338979.

Anonymous:1878:PCI

[335] Anonymous. The Permanent Commission of the International Statistical Congress. Journal of the Statistical Society, 41(3):549–550, September 1878. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338980.

Anonymous:1878:SCP

[336] Anonymous. Statistical Conferences at Paris. Journal of the Statistical Society, 41(3):550–551, September 1878. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338981.

Anonymous:1878:EFM

[337] Anonymous. Enteric fever and milk supply. Journal of the Statistical Society, 41 (3):551–559, September 1878. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338982.

Anonymous:1878:NNB

[338] Anonymous. Notices of new books. Journal of the Statistical Society, 41(3):

559–562, September 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338983.

Anonymous:1878:ALa

[339] Anonymous. Additions to the library.

Journal of the Statistical Society, 41(3): 562–564, September 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338984.

Anonymous:1878:ALb

[340] Anonymous. Additions to the library.

Journal of the Statistical Society, 41(3): 564–572, September 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338985.

Anonymous:1878:VIb

[341] Anonymous. Volume information. Journal of the Statistical Society, 41(4):689– 711, December 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339165.

Shaw-Lefevre:1878:OAG

[342] G. J. Shaw-Lefevre. The opening address of G. J. Shaw-Lefevre, Esq., M.P., President of the Statistical Society, delivered on Tuesday, the 19th November, 1878. Journal of the Statistical Society, 41(4):573–596, December 1878. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339166.

Jevons:1878:RSU

[343] W. Stanley Jevons. Remarks on the statistical use of the arithmometer. Journal of the Statistical Society, 41(4):597–601, December 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339167.

Ingram:1878:APS

[344] Professor Ingram. Address of the President of Section F of the British Association, at the forty-eighth meeting, held at Dublin, in August, 1878. Journal of the Statistical Society, 41(4):602–629, December 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339168.

Bateman:1878:SC

[345] Alfred Edmund Bateman. The statistics of Canada. Journal of the Statistical Society, 41(4):630–636, December 1878. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339169.

Price:1878:APD

[346] Bonamy Price. Address of the President of the department of economy and trade of the National Association for the Promotion of Social Science, at the twenty-second annual Congress held at Cheltenham, in October, 1878. Journal of the Statistical Society, 41(4):637–653, December 1878. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339170.

Anonymous:1878:AR

[347] Anonymous. Agricultural returns for 1878. Journal of the Statistical Society, 41(4):654–661, December 1878. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339171.

Anonymous:1878:CP

[348] Anonymous. The Commissariat of Paris. Journal of the Statistical Society, 41(4): 662–666, December 1878. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339172.

Anonymous:1878:NES

[349] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 41(4):666–672, December 1878. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339173.

Anonymous:1878:NAL

[350] Anonymous. Notes on additions to the library. Journal of the Statistical Society, 41(4):672–687, December 1878. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339174.

Anonymous:1879:VIa

[351] Anonymous. Volume information. Journal of the Statistical Society, 42(1):i– viii, March 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338986.

Mouat:1879:RFS

[352] Frederic J. Mouat. Reports on: 1. the fourth session of the Permanent Commission of the International Statistical Congress, held in Paris, 10th–20th July, 1878; 2. the first session of the International Congress of Demography and Medical Geography, held in Paris, 5th– 10th July, 1878; 3. the second session of the International Prison Congress, held in Stockholm, 20th–27th August, 1878. Journal of the Statistical Society, 42(1): 1–35, March 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338987.

Giffen:1879:FPC

[353] Robert Giffen. On the fall of prices of commodities in recent years. Journal of the Statistical Society, 42(1):36– 78, March 1879. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2338988.

Walford:1879:FWP

[354] Cornelius Walford. The famines of the world: Past and present — Part II. Journal of the Statistical Society, 42 (1):79–275, March 1879. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338989.

Anonymous:1879:FCH

[355] Anonymous. Financial and commercial history of 1878. Journal of the Statistical Society, 42(1):276–305, March 1879. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338990.

Anonymous:1879:PE

[356] Anonymous. Prices of exports from 1861 to 1877. Journal of the Statistical Society, 42(1):305–321, March 1879. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338991.

Anonymous:1879:FLD

[357] Anonymous. The fires of London during the year 1878, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 42(1):322–325, March 1879. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338992.

Anonymous:1879:EL

[358] Anonymous. English literature in 1878. Journal of the Statistical Society, 42(1): 325–328, March 1879. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338993.

Anonymous:1879:UBR

[359] Anonymous. University boat races and sun-spot cycles. Journal of the Statistical Society, 42(1):328–329, March 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338994.

Anonymous:1879:SCB

[360] Anonymous. A suggested coincidence between the death-rate and the motions of the planet Jupiter. Journal of the Statistical Society, 42(1):330–331, March 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338995.

Anonymous:1879:NESa

[361] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 42(1):331–334, March 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338996.

Anonymous:1879:NAL

[362] Anonymous. Notes on additions to the library. Journal of the Statistical Society, 42(1):334–344, March 1879. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338997.

Anonymous:1879:PR

[363] Anonymous. Periodical returns. Journal of the Statistical Society, 42(1):345– 368, March 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338998.

Hayter:1879:CVP

[364] Henry Heylyn Hayter. The colony of Victoria: Its progress and present position. Journal of the Statistical Society, 42(2):369–405, June 1879. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339143.

Bourne:1879:SPS

[365] Stephen Bourne. On some phases of the silver question. Journal of the Statistical Society, 42(2):406–455, June 1879. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339144.

Hancock:1879:FCE

[366] W. Neilson Hancock. The feasibility of compulsory education in Ireland. Journal of the Statistical Society, 42(2):456– 479, June 1879. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339145.

Anonymous:1879:GRC

[367] Anonymous. General results of the commercial and financial history of 1878. Journal of the Statistical Society, 42(2): 480–504, June 1879. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339146.

Anonymous:1879:BCH

[368] Anonymous. Bankers’ Clearing House returns. Journal of the Statistical Society, 42(2):504–505, June 1879. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339147.

Anonymous:1879:LSM

[369] Anonymous. Lloyd’s statistics of marine casualties for the year 1878. Journal of the Statistical Society, 42(2):505– 521, June 1879. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339148.

Anonymous:1879:SRL

[370] Anonymous. Statistical results of the last census of the population of France. Journal of the Statistical Society, 42(2): 522–528, June 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339149.

Vesselovsky:1879:MTB

[371] A. Vesselovsky. The movement of travellers between European Russia and foreign countries from 1857 to 1876. Journal of the Statistical Society, 42(2):528– 534, June 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339150.

Anonymous:1879:FLL

[372] Anonymous. Farmers’ and landowners’ losses. Journal of the Statistical Society, 42(2):535–538, June 1879. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339151.

Anonymous:1879:NESb

[373] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 42(2):539–543, June 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339152.

Anonymous:1879:NRA

[374] Anonymous. Notes on recent additions to the library. Journal of the Statistical Society, 42(2):543–558, June 1879. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339153.

Anonymous:1879:RCF

[375] Anonymous. Report of the Council for the financial year ended 31st December, 1878, and for the sessional year ending 30th June, 1879, presented at the fortyfifth anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, London, on the 30th of June, 1879, with an account of the proceedings at the meeting. Journal of the Statistical Society, 42(3):559–572, September 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339049.

Anonymous:1879:PFF

[376] Anonymous. Proceedings of the fortyfifth anniversary meeting. Journal of the Statistical Society, 42(3):573–578, September 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339050.

Ravenstein:1879:CLB

[377] E. G. Ravenstein. On the Celtic languages in the British Isles; a statistical survey. Journal of the Statistical Society, 42(3):579–643, September 1879. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339051.

Guy:1879:TA

[378] William A. Guy. On tabular analysis. Journal of the Statistical Society, 42 (3):644–662, September 1879. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339052.

Martin:1879:SEC

[379] John B. Martin. On some effects of a crisis on the banking interest. Journal of the Statistical Society, 42(3):663–708, September 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339053.

Anonymous:1879:GAP

[380] Anonymous. The gazette average prices of corn. Journal of the Statistical Society, 42(3):709–723, September 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339054.

Anonymous:1879:MSU

[381] Anonymous. Miscellaneous statistics of the United Kingdom. Journal of the Statistical Society, 42(3):724–727, September 1879. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339055.

Anonymous:1879:NESc

[382] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 42(3):727–736, September 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339056.

Anonymous:1879:ALN

[383] Anonymous. Additions to the library, with notes. Journal of the Statistical Society, 42(3):736–749, September 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339057.

Anonymous:1879:VIb

[384] Anonymous. Volume information. Journal of the Statistical Society, 42(4):883– 909, December 1879. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339081.

Brassey:1879:AEU

[385] Thomas Brassey. Agriculture in England and the United States. The inaugural address of Thomas Brassey, Esq., M.P., President of the Statistical Society, delivered on Tuesday, the 18th of November, 1879. Journal of the Statistical Society, 42(4):751–764, December 1879. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339082.

Anonymous:1879:PN

[386] Anonymous. Proceedings on the 18th November, 1879. Journal of the Statistical Society, 42(4):765–770, December 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339083.

Shaw-Lefevre:1879:APS

[387] G. Shaw-Lefevre. Address of the President of Section F, “Economic Science and Statistics,” of the British Association, at the forty-ninth meeting, held at Sheffield, in August, 1879. Journal of the Statistical Society, 42(4):771–783, December 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339084.

Anonymous:1879:SP

[388] Anonymous. Statistics in Parliament. Journal of the Statistical Society, 42(4): 784–790, December 1879. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339085.

Anonymous:1879:AR

[389] Anonymous. Agricultural returns for 1879. Journal of the Statistical Society, 42(4):790–808, December 1879. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339086.

Anonymous:1879:C

[390] Anonymous. The crops of 1879. Journal of the Statistical Society, 42(4):808–816, December 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339087.

Newmarch:1879:INB

[391] William Newmarch. The increase in the number of banks and branches in the Metropolis; the English counties; Scotland and Ireland; during the twenty years 1858–78, with deductions relating to the Clearing House returns, the increase on the average annual circulation of Bank of England. Journal of the Statistical Society, 42(4):817–844, December 1879. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339088.

Anonymous:1879:WMU

[392] Anonymous. With what margin of uncalled capital may the leading unlimited banks safely become limited? Journal of the Statistical Society, 42(4):844–848, December 1879. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339089.

Anonymous:1879:IPG

[393] Anonymous. The influence of the price of goods on the movements of the foreign trade of France. Journal of the Statistical Society, 42(4):848–857, December 1879. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339090.

Anonymous:1879:SAC

[394] Anonymous. Statistics of Australasian colonies. Journal of the Statistical Society, 42(4):858–859, December 1879. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339091.

Anonymous:1879:NSA

[395] Anonymous. Notes on some of the additions to the library. Journal of the Statistical Society, 42(4):860–873, December 1879. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339092.

Anonymous:1879:ALD

[396] Anonymous. Additions to the library during the quarter. Journal of the Statistical Society, 42(4):873–882, December 1879. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339093.